Some recent financial statements for Smolira Golf Corp. follow. 2017 and 2018 Balance Sheets Assets Liabilities and Owners Equity 2017 2018 2017 2018 Current assets Current liabilities Cash Accounts $39,216 Accounts payable $35,88539.216 29,116 406043,232 $ 39,37243,332 17,325 25,834 18,551 Notes payable Other 20,508 21,214 receivable Inventory Total $58,496111,564 81,09486,491 Total Long-term debt Owners equity $122,500190,733 Common stock and paid-in $ 56,500 $56,500 surplus Accumulated retained 264,332299,723 earnings Fixed assets Net plant and equipment $465,930$521,883 Total $320,832$356,223 Total liabilities and owners $524,426 $633,447 equity Total assets $524,426 $633,447 Prev 8of 1011 Next>
Co 2018 Income Statement Sales Cost of goods sold Depreciation $513,954 364,578 46,338 Earnings before interest and taxes $103,038 Interest paid 21,183 Taxable income Taxes (25%) $ 81,855 20,464 Net income $ 61,391 Dividends Retained earnings $26,000 35,391 Smolira Golf Corp. has 55,000 shares of common stock outstanding, and the market price for a share of stock at the end of 2018 was $32. a. What is the price-earnings ratio? (Do not round intermediate calculations and b. What are the dividends per share? (Round your answer to 2 decimal places, e.g. c. What is the market-to-book ratio at the end of 2018? (Round your answer to 2 d. If the companys growth rate is 8 percent, what is the PEG ratio? (Do not round round your answer to 2 decimal places, e.g., 32.16.) 32.16.) decimal places, e.g.. 32.16.) intermediate calculations and round your answer to 2 decimal places, e.g, 32.16.)
ee Toune your answer to 2 a. Price-earnings ratio b. Dividends per share c, | Market-to-book ratio d. PEG ratio imes imes times K Pre
Answer #1

(A). PRICE EARNING RATIO = MARKET PRICE PER SHARE (MPS) / EARNING PER SHARE (EPS)

$32 / $1.1162

=28.67

(EPS = NET INCOME/ NO. OF SHARES) , (MPS IS GIVEN)

(EPS = 61391 / 55000)

(B) DIVIDEND PER SHARE = DIVIDEND / NO. OF SHARES

= $26000 / 55000

= $0.47

(C) MARKET TO BOOK RATIO = MARKET VALUE OF FIRM / NET ASSETS OF FIRM

= $32*55000 / $(56500 + 299723)

= 1760000 / 356223

= 4.94

(NET ASSETS HAS BEEN CALCULATED BY ADDING OWNER'S EQUITY AND RETAINED EARNINGS. ALTERNATIVELY, IT CAN ALSO BE CALCULATED BY SUBTRACTING ALL THE OUTSIDE LIABILITIES FROM TOTAL ASSETS)

(D). PEG RATIO = P.E. RATIO / GROWTH RATE

= 28.67 / 8

= 3.58